Write the polynomial equation m^6 + 5m^3 - 10 = 0 in quadratic form.
How do I do that^^? Please Help!

Answer by fcabanski(1391)

Quadratic form means the highest order (power) is 2 (such as x^2).


m^6 = (m^3)^2. So make u= m^3 and rewrite the equation with u.


u^2 + 5u - 10 = 0
Check my profile and contact me for one on one online tutoring. If you need more help with this topic, a live session will make it clear.